# Height filtrations and base loci on flag bundles over a curve

### v1

@@ 3. Reinforcement of Zhang’s inequality into an equality
Let $G=\operatorname{GL}_{4}$, $E$ be a vector bundle on $C$ of rank $4$ and consider the $\operatorname{Gr}(2,4)$-bundle $\operatorname{Gr}_{2}(E)$. In this case the string polytopes (with respect to a certain reduced decomposition) is just the Gelfand-Zetlin polytope [18, Remark 2.4]. Identify $X(T)_{\mathbb{Q}}$ with $\bigoplus_{i}\mathbb{Q}\lambda_{i}$ as before. The line bundle $\mathcal{O}(1)$ correspond to the character $\lambda=(0,0,1,1)$.
@@
We see that $\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ is also $\frac{1}{2}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4})$.
@@ 4. Base loci and movable cones
4.1. Preliminaries on algebraic geometry
We briefly recall some general theory of base loci. Let $\mathbf{k}$ be a field. In our case, it would be either the base field $k$ or the function field $K=K(C)$. Let $Y\rightarrow\operatorname{Spec}\mathbf{k}$ be a normal projective variety over $\mathbf{k}$, and $L$ be a line bundle over $X$. Let $V$ be a $\mathbf{k}$-subspace of $H^{0}(Y,L)$. Consider the evaluation map
$H^{0}(Y,L)\otimes_{\mathbf{k}}L^{\vee}\rightarrow\mathcal{O}_{Y}.$
The base locus $\mathrm{Bs}(V)$ of $V$ is the reduced scheme associated to the closed subscheme defined by the ideal sheaf given by the image of $\pi^{*}V\otimes{L^{\vee}}\rightarrow\mathcal{O}_{Y}$.
Definition 4.1.
The stable base locus of $L$ is defined as $\mathrm{B}(L)=\bigcap_{n\geq 1}\mathrm{Bs}(H^{0}(Y,nL))$. Take any ample line bundle $A$. The augmented base locus of $L$ is defined as $\mathrm{B}_{+}(L)=\bigcap_{n\geq 1}\mathrm{B}(nL-A)$, which is independent of the choice of $A$.
The following characterization of augmented base loci in terms of restricted volumes is well known, which we include here for the reader’s convenience.
@@ 4.1. Preliminaries on algebraic geometry
Let $Z$ be a closed subvariety of $Y$. The restricted volume is defined as
$\operatorname{Vol}(Y|Z,L)=\limsup_{n\rightarrow+\infty}\frac{\dim_{\mathbf{k}}\mathrm{Im}(H^{0}(Y,nL)\rightarrow H^{0}(Z,nL|_{Z}))}{n^{\dim Z}/(\dim Z)!}$
where Im()normal-$\mathrm{Im}(\cdot)$ denotes the image. Then $\mathrm{B}_{+}(L)=\bigcup_{\operatorname{Vol}(Y|Z,L)=0}Z$.
Let $\mathrm{Pic}(Y)$ be the picard group of $Y$. Since $\mathrm{B}(L)=\mathrm{B}(mL)$ for any $m>1$, we can extend the definitions of $\mathrm{B}(\cdot)$ and $\mathrm{B}_{+}(\cdot)$ to $\mathrm{Pic}(Y)_{\mathbb{Q}}:=\mathrm{Pic}(Y)\otimes\mathbb{Q}$. Moreover, let $N^{1}(Y):=\mathrm{Div}(X)/\equiv_{num}$ be the numerical group of divisors. It is proved in [13] that the augmented base loci are well defined on $N^{1}(Y)_{\mathbb{R}}:=N^{1}(Y)\otimes_{\mathbb{Z}}\mathbb{R}$.
Let $\mathcal{Y}$ be a projective variety with a surjective morphism $\mathcal{Y}\rightarrow C$, and $\mathcal{L}$ be a line bundle over $\mathcal{Y}$. For any scheme-theoretical point $p\in C$, we denote by $\kappa(p)$ the residue field of $p$. Let $\mathcal{Z}$ be a subvariety of $\mathcal{Y}$. The fiber $\mathcal{Z}\times_{C}\mathrm{Spec}(\kappa(p))$ of $\mathcal{Z}$ over $p$, is denoted as $\mathcal{Z}_{p}$. We also denote by $\mathcal{L}_{p}$ the restriction $\mathcal{L}|_{\mathcal{Y}_{p}}$ of $\mathcal{L}$ to $\mathcal{Y}_{p}$.
We have the following lemma, which will be useful for our computation of augmented base loci in the next subsection.
@@
The restriction of $\mathrm{B}(\mathcal{L})$ to $\mathcal{Y}_{p}$ can be computed by
$B(\mathcal{L})_{p}=\bigcap\nolimits_{n\geq 1}\mathrm{Bs}(\mathrm{Im}(H^{0}(\mathcal{X},n\mathcal{L})\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{Y}_{p},n\mathcal{L}_{p}))).$
4.2. Augmented base loci on the flag bundle $\mathcal{X}$
Let $\lambda\in X(P)$ be a strictly anti-dominant character. We fix a closed point $p_{0}\in C$, and set $f=\pi^{*}(\mathcal{O}_{C}(p_{0}))$. In this subsection, we are going to compute the augmented base locus $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$. Remind that we have the decomposition
$H^{0}(G/P,M_{\lambda})=\bigoplus_{\nu\in X(T)}H^{0}(G/P,M_{\lambda})[\nu]$
@@ 4.2. Augmented base loci on the flag bundle $\mathcal{X}$
$H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\geq t}H^{0}(G/P,M_{\lambda})[\nu].$
We start with a lemma describing the vanishing behaviour of sections in $H^{0}(G/P,M_{\lambda})[\nu]$.
@@
    $s\in H^{0}(G/P,M_{\lambda})[\nu]$ for $\nu\not\leq w\lambda$Longrightarrow $s(x)=0$ for any $x\in\overline{BwP}/P$.
@@
As for $(2)$, for any $\nu\not\leq w\lambda$, $H^{0}(\overline{BwP}/P,M_{\lambda})[\nu]=0$. Consequently $\mathrm{Res}(s)=0$ for any $s\in H^{0}(G/P,M_{\lambda})[\nu]$.
@@
Let $p$ be a scheme-theoretical point of $C$. Consider the filtration $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})$ given by
$H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}=F_{Q,\kappa(p)}\times_{Q}H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}$
for $t\in\mathbb{R}$, we have $\operatorname{Bs}\Big{(}H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}\Big{)}=\mathcal{X}_{p}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}\mathcal{C}_{w,p}$.
@@
Let $F_{T,\kappa(p)}$ be a reduction of $F_{Q,\kappa(p)}$ to $T$. For any $\nu\in X(T)$, consider the $\kappa(p)$-subspace $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\nu]=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})[\nu]$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})$. For $w^{\prime}\in W/W_{P}$, set $C_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}Bw^{\prime}P/P$ and $X_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}\overline{Bw^{\prime}P}/P$. Then $\mathcal{C}_{w,p}=\coprod_{\pi(w^{\prime})=w}C_{w^{\prime}}$ and $\mathcal{X}_{w,p}=\bigcup_{\pi(w^{\prime})=w}X_{w^{\prime}}$ where $\pi:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ is the natural projection.
@@
On the other hand, we claim that if $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$, then for any $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}$, the restriction of $s$ to $X_{w}$ is zero. In fact, we may assume $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\nu]$ for some $\nu$ and $\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\geq t$ implies that $\nu\not\leq w^{\prime}\lambda$ for any $w^{\prime}\in\pi^{-1}(w)$ (because if $\nu\leq w^{\prime}\lambda$ for some $w^{\prime}$, then $\langle{\mathrm{deg}}(F_{Q}),\nu\rangle\leq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$). By Lemma 4.4(2), the restriction of $s$ to $X_{w^{\prime}}$ is zero for each $w^{\prime}$. We win. ∎
Now we are ready to compute the augmented base locus of $\mathcal{L}_{\lambda}-tf$.
@@
The augmented base locus of $\mathcal{L}_{\lambda}-tf$ is given by
@@
We may assume that $t\in\mathbb{N}$ by taking a sufficiently large tensor power of $\mathcal{L}_{\lambda}-tf$. Let $t_{0}=\min\{\langle{\mathrm{deg}}(F_{Q}),\nu\rangle:H^{0}(G/P,M_{\lambda})[\nu]\not=0\text{ and }\langle{\mathrm{deg}}(F_{Q}),\nu\rangle>t\}$. Take $E=F_{Q}\times_{Q}H^{0}(G/P,M_{\lambda})_{t_{0},{\mathrm{deg}}(F_{Q})}\otimes\mathcal{O}_{C}(-tx)$, which is a subbundle of $\pi_{*}(\mathcal{L}_{\lambda}-tf).$ Moreover, $\mu_{\min}(E)=t_{0}-t>0$, i.e. $E$ is ample.
@@
Now we take an ample line bundle $\mathcal{A}$ over $\mathcal{X}$ of form $\mathcal{L}_{\lambda^{\prime}}-t^{\prime}f$ where $\lambda^{\prime}\in X(T)$ is a strictly anti-dominant character and $\langle{\mathrm{deg}}(F_{Q}),\nu\rangle-t^{\prime}>0$ for any $\nu\in X(T)$ such that $H^{0}(G/P,M_{\lambda^{\prime}})[\nu]\not=0.$
@@
we have $\operatorname{Vol}(\mathcal{L}_{\lambda}-tf|_{\mathcal{X}_{w}})=0$ by a comparison result [12, Theorem 2.4] deduced from the Riemann-Roch theorem. Therefore $\operatorname{Vol}(\mathcal{X}|\mathcal{X}_{w},\mathcal{L}_{\lambda}-tf)=0$, which implies that $\mathcal{X}_{w}\subseteq B_{+}(\mathcal{L}_{\lambda}-tf)$ due to Theorem 4.2. In conclusion, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
Remark 4.7.
Comparing Theorem 2.11 and Theorem 4.6, we find that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)_{K}$ agrees with the height filtration $Z_{t}$ for every $t\not=\zeta_{w}$. At critical points, $\mathrm{B}_{+}$ is upper continuous while $Z_{t}$ is lower continuous.
4.3. Positive cones of flag bundles over curves
We first give a reminder on the definitions of various positive cones of a projective variety. Let $Y$ be a smooth projective variety over a field $\mathbf{k}$ (of characteristic zero). We recall the following:
  1. (1)
@@ 4.3. Positive cones of flag bundles over curves
  3. (3)
    For each $k=1,\cdots,\dim Y$, the $k$-th movable cone is defined as $\operatorname{Mov}^{k}(Y)=\big{\{}D\in N^{1}(Y)_{\mathbb{R}}:\mathrm{codim}(\mathrm{B}_{+}(D))\geq k\big{\}}$ which is open in $N^{1}(Y)_{\mathbb{R}}$. They are open cones in $N^{1}(Y)_{\mathbb{R}}$ and we have obvious inclusions $\operatorname{Mov}^{d}\subseteq\cdots\subseteq\operatorname{Mov}^{1}$. Note that $\operatorname{Mov}^{1}$ is the big cone and $\operatorname{Mov}^{d}$ is the ample cone.
Now we come back to the case of flag varieties. Denote by
@@
Since $G$ and $P$ are connected linear algebraic groups over a characteristic zero field, both $\mathrm{Pic}(G)$ and $\operatorname{Pic}(P)$ are finite. Notice that the numerical equivalence and the linear equivalence coincide on $G/P$ since $G/P$ is smooth fano. Therefore the linear map $X(P)_{\mathbb{R}}\rightarrow N^{1}(G/P)_{\mathbb{R}}$ is surjective. It is well-known that $M_{\lambda}$ is ample $\Longleftrightarrow$ $M_{\lambda}$ is big $\Longleftrightarrow$ $\lambda$ is strictly anti-dominant, thus
@@
By the above description, we define two functions $\langle\alpha^{\vee},\cdot\rangle$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ on $N^{1}(F/P)_{\mathbb{R}}$ as
@@
Corollary 4.10.
$\mathcal{L}_{\lambda}-tf\in\operatorname{Mov}^{k}(F/P)$ normal-$\Longleftrightarrow$ $e_{k}(h_{\mathcal{L}_{\lambda}})>t$ where $e_{k}$ is the $k$-th minimum of Zhang.

### v2

@@ 3.2. String polytopes and weight polytopes
Let $G=\operatorname{GL}_{4}$, $E$ be a vector bundle on $C$ of rank $4$ and consider the $\operatorname{Gr}(2,4)$-bundle $\operatorname{Gr}_{2}(E)$. In this case the string polytopes (with respect to a certain reduced decomposition) is just the Gelfand-Zetlin polytope [15, Remark 2.4]. Identify $X(T)_{\mathbb{Q}}$ with $\bigoplus_{i}\mathbb{Q}\lambda_{i}$ as before. The line bundle $\mathcal{O}(1)$ corresponds to the character $\lambda=(0,0,1,1)$.
@@
Thus $\frac{1}{\#W/W_{P}}\sum_{w\in W/W_{P}}\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle=\frac{1}{2}(\mu_{1}+\mu_{2}+\mu_{3}+\mu_{4})=h(X)$, which is predicted by Theorem 3.2.
@@ 4. Base loci and movable cones
Keep the notations in §2. We fix a closed point $p_{0}\in C$ and set $f=\pi^{*}(\mathcal{O}_{C}(p_{0}))$. For any $w\in W_{Q}\backslash W/W_{P}$, let $\mathcal{C}_{w}=F_{Q}\times_{Q}QwP/P$. Let $t\in\mathbb{Q}$. The main theorem of this section is a computation of base loci.
Theorem 4.1.
Let $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ be the augmented base locus of $\mathcal{L}_{\lambda}-tf$. Then
Comparing Theorem 2.1 and Theorem 4.1, we find that $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ agrees with the Zariski closure of $Z_{t}$ in $\mathcal{X}$ for every $t\not=\zeta_{w}$. At critical points, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)$ is upper continuous while $Z_{t}$ is lower continuous.
As a corollary, we compute the movable cones on flag bundles. (Please see §4.3 for the notations.)
Theorem 4.2.
The $k$-th movable cone $\operatorname{Mov}^{k}(F/P)$ is the cone defined by
  1. (1)
  2. (2)
In particular, let $\mathcal{L}$ be a line bundle on $F/P$ and $e_{k}(h_{\mathcal{L}})$ be the $k$-th minimum of Zhang of the height function $h_{\mathcal{L}}$, we have
Corollary 4.3.
$\mathcal{L}$ is $k$-movable if and only if $e_{k}(h_{\mathcal{L}})>0$.
4.1. Augmented base loci and the restricted volume
Let $\mathbf{k}$ be a field and $Y$ be a normal projective variety over $\mathbf{k}$. Let $L$ be a line bundle on $Y$ and $V\subseteq H^{0}(Y,L)$ be a subspace. We shall denote by $\mathrm{Bs}(V)$ the base locus of $V$.
The stable base locus of $L$ is defined as $\mathrm{B}(L)=\bigcap_{n\geq 1}\mathrm{Bs}(H^{0}(Y,nL))$. The augmented base locus of $L$ is defined as $\mathrm{B}_{+}(L)=\bigcap_{n\geq 1}\mathrm{B}(nL-A)$ for any ample line bundle $A$. This is independent of the choice of $A$.
Let $Z$ be a closed subvariety of $Y$ of dimension $d$. The restricted volume is defined as
$\operatorname{Vol}(Y|Z,L)=\limsup_{n\rightarrow\infty}\frac{\dim_{\mathbf{k}}\mathrm{Im}\Big{(}H^{0}(Y,nL)\rightarrow H^{0}(Z,nL|_{Z})\Big{)}}{n^{d}/d!}.$
The following characterization of augmented base loci in terms of restricted volumes is well known, which we include here for readers’ convenience.
@@ 4.1. Augmented base loci and the restricted volume
$\mathrm{B}_{+}(L)$ is the union of all closed subvarieties $Z$ with $\operatorname{Vol}(Y|Z,L)=0$.
Let $N^{1}(Y)$ be the Néron-Severi group of $Y$. It is proved in [10] that augmented base loci are well defined on $N^{1}(Y)_{\mathbb{R}}:=N^{1}(Y)\otimes_{\mathbb{Z}}\mathbb{R}$.
Let $\mathcal{Y}\longrightarrow C$ be a projective flat family over the curve $C$, and $\mathcal{L}$ be a line bundle over $\mathcal{Y}$. For any scheme theoretical point $p\in C$, we denote by $\kappa(p)$ the residue field of $p$. Let $\mathcal{Z}$ be a closed subvariety of $\mathcal{Y}$, we denote by $\mathcal{Z}_{p}$ the fiber $\mathcal{Z}\times_{C}\mathrm{Spec}(\kappa(p))$ of $\mathcal{Z}$ over $p$. We denote by $\mathcal{L}_{p}$ the restriction $\mathcal{L}|_{\mathcal{Y}_{p}}$ of $\mathcal{L}$ to $\mathcal{Y}_{p}$.
@@
$\mathrm{B}(\mathcal{L})_{p}=\bigcap\nolimits_{n\geq 1}\mathrm{Bs}\left(\mathrm{Im}\Big{(}H^{0}(\mathcal{Y},n\mathcal{L})\otimes_{k}\kappa(p)\rightarrow H^{0}(\mathcal{Y}_{p},n\mathcal{L}_{p})\Big{)}\right)$.
4.2. Proof of Theorem 4.1
Let $M_{\lambda}=G\times_{P}k_{\lambda}$, which is an ample line bundle on $G/P$. Recall that we have the decomposition
$H^{0}(G/P,M_{\lambda})=\bigoplus_{\mu\in X(T)}H^{0}(G/P,M_{\lambda})[\mu]$
@@
$H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}=\bigoplus_{\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\geq t}H^{0}(G/P,M_{\lambda})[\mu].$
We start with a lemma describing the vanishing behaviour of sections in $H^{0}(G/P,M_{\lambda})[\mu]$.
@@
    $s\in H^{0}(G/P,M_{\lambda})[\mu]$ for $\mu\not\leq w\lambda$Longrightarrow $s(x)=0$ for any $x\in\overline{BwP}/P$.
@@
As for $(2)$, for any $\mu\not\leq w\lambda$, $H^{0}(\overline{BwP}/P,M_{\lambda})[\mu]=0$. Consequently, $\mathrm{Res}(s)=0$ for any $s\in H^{0}(G/P,M_{\lambda})[\mu]$. ∎
Let $p$ be a scheme theoretical point of $C$ and consider the filtration $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{\bullet,{\mathrm{deg}}(F_{Q})}$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})$ given by
$H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}=F_{Q,\kappa(p)}\times_{Q}H^{0}(G/P,M_{\lambda})_{t,{\mathrm{deg}}(F_{Q})}.$
We have
@@
$\operatorname{Bs}\Big{(}H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}\Big{)}=\mathcal{X}_{p}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle\geq t}\mathcal{C}_{w,p}$.
@@
Let $F_{T,\kappa(p)}$ be a reduction of $F_{Q,\kappa(p)}$ to $T$. For any $\mu\in X(T)$, consider the $\kappa(p)$-subspace $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\mu]=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})[\mu]$ of $H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})=F_{T,\kappa(p)}\times_{T}H^{0}(G/P,M_{\lambda})$. For $w^{\prime}\in W/W_{P}$, set $C_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}Bw^{\prime}P/P$ and $X_{w^{\prime}}=F_{T,\kappa(p)}\times_{T}\overline{Bw^{\prime}P}/P$. Then $\mathcal{C}_{w,p}=\coprod_{\pi(w^{\prime})=w}C_{w^{\prime}}$ and $\mathcal{X}_{w,p}=\bigcup_{\pi(w^{\prime})=w}X_{w^{\prime}}$ where $\pi:W/W_{P}\longrightarrow W_{Q}\backslash W/W_{P}$ is the natural projection.
@@
On the other hand, we claim that if $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$, then for any $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})_{t,{\mathrm{deg}}(F_{Q})}$, the restriction of $s$ to $X_{w}$ is zero. In fact, we may assume $s\in H^{0}(\mathcal{X}_{p},\mathcal{L}_{\lambda,p})[\mu]$ for some $\mu$ and $\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\geq t$ implies that $\mu\not\leq w^{\prime}\lambda$ for any $w^{\prime}\in\pi^{-1}(w)$ (because if $\mu\leq w^{\prime}\lambda$ for some $w^{\prime}$, then $\langle{\mathrm{deg}}(F_{Q}),\mu\rangle\leq\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle<t$). By Lemma 4.6(2), the restriction of $s$ to $X_{w^{\prime}}$ is zero for each $w^{\prime}$. We win. ∎
@@
Let $B_{+}(\mathcal{L}_{\lambda}-tf)$ be the augmented base locus of $\mathcal{L}_{\lambda}-tf$. Then
@@
We may assume that $t\in\mathbb{N}$ by taking a sufficiently large tensor power of $\mathcal{L}_{\lambda}-tf$. Let $t_{0}=\min\{\langle{\mathrm{deg}}(F_{Q}),\mu\rangle:H^{0}(G/P,M_{\lambda})[\mu]\not=0\text{ and }\langle{\mathrm{deg}}(F_{Q}),\mu\rangle>t\}$. Take $E=F_{Q}\times_{Q}H^{0}(G/P,M_{\lambda})_{t_{0},{\mathrm{deg}}(F_{Q})}\otimes\mathcal{O}_{C}(-tx)$, which is a subbundle of $\pi_{*}(\mathcal{L}_{\lambda}-tf).$ Moreover, $\mu_{\min}(E)=t_{0}-t>0$, i.e. $E$ is ample.
@@
Now we take an ample line bundle $\mathcal{A}$ over $\mathcal{X}$ of form $\mathcal{L}_{\lambda^{\prime}}-t^{\prime}f$ where $\lambda^{\prime}\in X(T)$ is a strictly anti-dominant character and $\langle{\mathrm{deg}}(F_{Q}),\mu\rangle-t^{\prime}>0$ for any $\mu\in X(T)$ such that $H^{0}(G/P,M_{\lambda^{\prime}})[\mu]\not=0.$
@@
we have $\operatorname{Vol}(\mathcal{L}_{\lambda}-tf|_{\mathcal{X}_{w}})=0$ by a comparison result [9, Theorem 2.4] deduced from the Riemann-Roch theorem. Therefore $\operatorname{Vol}(\mathcal{X}|\mathcal{X}_{w},\mathcal{L}_{\lambda}-tf)=0$, which implies that $\mathcal{X}_{w}\subseteq B_{+}(\mathcal{L}_{\lambda}-tf)$ due to Theorem 4.4. In conclusion, $\mathrm{B}_{+}(\mathcal{L}_{\lambda}-tf)=\mathcal{X}\backslash\coprod_{\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle>t}\mathcal{C}_{w}.$
4.3. Movable cones and proof of Theorem 4.2
Let $Y$ be a projective smooth variety over $\mathbf{k}$. Recall
@@
  • For each $k=1,\cdots,\dim Y$, the $k$-th movable cone is defined as $\operatorname{Mov}^{k}(Y)=\big{\{}D\in N^{1}(Y)_{\mathbb{R}}:\mathrm{codim}(\mathrm{B}_{+}(D))\geq k\big{\}}$ which is a open in $N^{1}(Y)_{\mathbb{R}}$.
We have obvious inclusions $\operatorname{Mov}^{d}(Y)\subseteq\cdots\subseteq\operatorname{Mov}^{1}(Y)$. Note that $\operatorname{Mov}^{1}(Y)$ is just the big cone and $\operatorname{Mov}^{d}(Y)$ is just the ample cone.
Now we go back to the setting of §2. Denote by
@@
Since $G$ and $P$ are connected linear algebraic groups over a characteristic zero field, both $\mathrm{Pic}(G)$ and $\operatorname{Pic}(P)$ are finite. Notice that the numerical equivalence and the linear equivalence coincide on $G/P$ since it is smooth and Fano. Therefore the linear map $X(P)_{\mathbb{R}}\rightarrow N^{1}(G/P)_{\mathbb{R}}$ is surjective. It is well-known that $M_{\lambda}$ is ample $\Longleftrightarrow$ $M_{\lambda}$ is big $\Longleftrightarrow$ $\lambda$ is strictly anti-dominant, thus
@@
Thanks to the above description, we may define two functions $\langle\alpha^{\vee},\cdot\rangle$ and $\langle{\mathrm{deg}}(F_{Q}),w\cdot\rangle$ on $N^{1}(F/P)_{\mathbb{R}}$ as